$SPAD/src/input f01bsf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
n := 6
nz := 15
nzmax := 50
licn := 3*nzmax
ta1:List SF := new(licn-nz,0.0);
ta2:List SF := [5.,2.,-1.,2.,3.,-2.,1.,1.,-1.,-1.,2.,-3.,-1.,-1.,6.] ;
ta3 := concat(ta2,ta1);
a :Matrix SF := matrix [ta3]
lirn := (3*nzmax/2)::Integer
tr1:List Integer := new(lirn-nz,0);
tr2 := [1,2,2,2,3,4,4,4,5,5,5,5,6,6,6];
tr3 := concat(tr2,tr1);
irn := matrix[tr3]
tc1:List Integer := new(licn-nz,0);
tc2 := [1,2,3,4,3,1,4,5,1,4,5,6,1,2,6];
tc3 := concat(tc2,tc1);
icn :=  matrix [tc3]
pivot :SF:= 0.1
lblock := true
grow := true
abort := [true,true,false,true]
ifail := 110
f01brfResult:=f01brf(n,nz,licn,lirn,pivot,lblock,grow,abort,a,irn,icn,ifail);
eta :SF:= .1 -- now run f01bsf
abort := true
keep := f01brfResult.'ikeep;
disp := f01brfResult.'idisp;
cn :=  f01brfResult.'icn;
ta2:=[10.,12.,-3.,-1.,15.,-2.,10.,-1.,-1.,-5.,1.,-1.,-1.,-2.,6.];
a:= matrix [concat (ta2,ta1)]
result:=f01bsf(n,nz,licn,matrix [tr2],matrix [tc2],cn,keep,grow,eta,abort,_
matrix [[disp(1,1),disp(1,2)]],a,ifail)

———-

References^{[1] nothing}

4